OpenStudy (marinadanyel):

What is the graph of the function f(x) = the quantity of x squared plus 3 x minus 4, all over x plus 4 ?

OpenStudy (marinadanyel):

\[\frac{ x^2 + 3x - 4 }{ x + 4 }\]

jimthompson5910 (jim_thompson5910):

hint: try factoring the numerator

OpenStudy (photon336):

heh, or we could factor out the top and simplify. \[\frac{ (x+4)(x-1) }{ (x+4) }\]

OpenStudy (marinadanyel):

What would my Y-int be?

OpenStudy (marinadanyel):

-4?

jimthompson5910 (jim_thompson5910):

Look at @Photon336 's last post. Notice that the numerator is factored. A pair of terms will cancel.

jimthompson5910 (jim_thompson5910):

what is left behind?

OpenStudy (marinadanyel):

Oh, so x-1

jimthompson5910 (jim_thompson5910):

yes, so that complicated rational expression is equivalent to x-1 The only subtle difference is that there's a hole on the line y = x-1. Where is the hole?

OpenStudy (photon336):

yeah heh, always see if you can factor the numerator or denominator in these problems, to see if you can simplify the question any further.

OpenStudy (photon336):

@marinadanyel if you need to graph functions desmos is a good site to do that.

OpenStudy (marinadanyel):

Where does the Y-int come in?

OpenStudy (marinadanyel):

I'm not sure where the hole is

OpenStudy (trojanpoem):

At x = -4 , the zeros of the denominator

jimthompson5910 (jim_thompson5910):

Think of y = x - 1 as y = 1*x + (-1) The equation `y = 1*x + (-1)` is in the form `y = mx+b`

OpenStudy (marinadanyel):

(3, -7)

OpenStudy (marinadanyel):

that's where the hole is?

jimthompson5910 (jim_thompson5910):

close but no

jimthompson5910 (jim_thompson5910):

notice how -4 makes the denominator of the original expression equal to 0 x+4 = -4+4 = 0 so the hole will be when x = -4 plug x = -4 into y = x-1 to find the y coordinate of this hole

OpenStudy (marinadanyel):

(-4, -5)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (marinadanyel):

Okay, I see, I didn't understand the substitution at first.

OpenStudy (marinadanyel):

Thanks!

jimthompson5910 (jim_thompson5910):

no problem